OpenStudy (anonymous):

How do you multiply using absolute value like 14-2|-4-2(1.5)|

OpenStudy (anonymous):

you look at what is inside of the absolute value first. -4-2*1.5=-7. The absolute value of -7=7. Now we have our normal arithmetic.14-2*7=14-14=0

OpenStudy (anonymous):

Ok..the answer to the problem is 4.

OpenStudy (anonymous):

I know the answer I just can't figure out how to get it.